OpenStudy (anonymous):

Which equations in point-slope form are equations of the line that pass through the points (4, 5) and (−3, −1)? Choose exactly two answers that are correct.

OpenStudy (anonymous):

A: y - 3 = 7/6 (x - 1) B: y - 5 = 6/7 (x - 4) C: y + 1 = 6/7 (x + 3) D: y - 4 = 7/6 (x - 5)

OpenStudy (solomonzelman):

find the slope first.

OpenStudy (solomonzelman):

\(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2}}{\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2}} }\) where \(\LARGE \color{black}{ \displaystyle {\rm m} }\) is the slope \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\) are your two points.

OpenStudy (solomonzelman):

plase find the slope for me...

OpenStudy (anonymous):

7

OpenStudy (solomonzelman):

no

OpenStudy (anonymous):

6

OpenStudy (solomonzelman):

\(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{{\rm 5}}-\color{red}{{\rm (-1)}}}{\color{green}{{\rm 4}}-\color{darkgoldenrod}{{\rm (-3)}}} }\)

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

can you finish finding the slope from there?

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

4

OpenStudy (solomonzelman):

\(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{{\rm 5}}-\color{red}{{\rm (-1)}}}{\color{green}{{\rm 4}}-\color{darkgoldenrod}{{\rm (-3)}}} }\) \(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{{\rm 5}}+\color{red}{{\rm 1}}}{\color{green}{{\rm 4}}+\color{darkgoldenrod}{{\rm 3}}} =? }\)

OpenStudy (anonymous):

it's either 13 or 7/6

OpenStudy (solomonzelman):

which one is it?

OpenStudy (anonymous):

7/6

OpenStudy (solomonzelman):

none of the above you posted.

OpenStudy (anonymous):

Now I'm confused

OpenStudy (solomonzelman):

can you add on top and bottom ?

OpenStudy (solomonzelman):

5+1 = ? 4+3= ? then re-write the fraction

OpenStudy (anonymous):

6/7

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

now, you need to find the line.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y-y_{_1}=m(x-x_{_1}) }\)

OpenStudy (solomonzelman):

use this formula. \(\large\color{black}{ \displaystyle (x_{_1},~~y_{_1}) }\) is the point \(\large\color{black}{ \displaystyle m }\) is the slope

OpenStudy (solomonzelman):

you can use it with any of the given to you points.

OpenStudy (anonymous):

(x - 4) and (x + 3) are the points i got